Using laws of exponents, simplify and write the answer in exponential form: 3^2 × 3^4 × 3^8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of dividing the powers with same base = a^m x a^n= a^m + n
Then,
= (3)^2 + 4 + 8
= 3^14

Thanks